# Efficient HDR Reconstruction from Real-World Raw Images

### v5

@@ 5.2 Comparison with the Other Methods
Evaluation on HDR sensor dataset. To validate the validity of our method on the real-world HDR sensor dataset, we utilize the Chen [4] test dataset for cross-validation, which has raw images captured by a Sony IMX267 image sensor. Compared with previous methods, our method achieves state-of-the-art performance in visual quality and quantitative metrics. The visual results from tests on the HDR sensor raw dataset (as shown in Fig. 6) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from severe color casts. Furthermore, results in Tab. 3 reveal that compared to AHDRNet [34], our method yields an improvement of more than 0.35 dB and 0.05 gain in PSNR and $\Delta E$, respectively, for scenes with an exposure ratio of 8. On average, our method attains gains exceeding 0.1 dB and 0.15 in PSNR and $\Delta E$.
@@
Table 6: Performance comparison of different HDR models on RealRaw-HDR dataset with realistic exposure ratios. The FLOPs are measured on the raw image of $7808\times 5824$ resolution. The best and second-best performances are in bold and underlined, respectively.
@@
Evaluation on FUJI raw dataset. We then evaluate our method on the FUJI raw datasets, which are real-world bracketed exposure raw images captured by the FUJI-FILM GFX50S II camera. Fig. 7 compares results from two high dynamic range scenes, where our method achieves significantly better visualization. Our method can recover both fine details in overexposed regions and rich colors in underexposed areas without introducing artifacts (see rows 1 and 2). Compared to AHDRNet, our method can effectively remove noise and preserve the structure of dark regions. Notably, the alignment module in DeepHDR, AHDRNet, HDR-Transformer, and NHDRRNet requires many line buffers, making it challenging to deploy on resource-limited edge devices. HDR-Transformer fails to perform inference even on RTX 3090 devices. In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images, as evident in the visual results of Fig. 4, Fig. 6, and Fig. 7.
@@
5.3 Ablation Study
This section investigates the raw LDR-HDR pair formation pipeline and the importance of different components in the whole RepUNet. We ablate the baseline model step by step and compare the performance differences.
Generalization of our LDR-HDR pair formation pipeline. Our raw LDR-HDR pair formation pipeline is proposed to generate paired raw LDR-HDR data but can also be adapted to generate paired sRGB HDR data. To demonstrate such generalization, we transform the collected RealRaw-HDR dataset with a fixed ISP pipeline into the sRGB color space, named the Raw2RGB-HDR dataset. For comparison, we train the sRGB HDR method AHDRNet [34] on our Raw2RGB-HDR dataset and Kalantari dataset [14] (taking the first two exposures as input, 74 pairs of images), respectively. The test dataset is from the Kalantari dataset. Results in Tab. 7 show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on the Kalantari dataset by 2.86 dB in PSNR. The performance gains benefit from an efficient and user-friendly data acquisition pipeline that generates more trainable data pairs. The results demonstrate that our data pipeline is also effective in generating paired LDR-HDR data in sRGB space.
Table 7: We train the sRGB HDR method AHDRNet on our Raw2RGB-HDR dataset and Kalantari dataset, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR dataset.
@@ 5.3 Ablation Study
Loss functions. To test the effects of alignment-free and motion-aware short-exposure-first selection loss and Bayer loss, we set the $L2$ joint $L_{\textrm{ssim}}$ loss as the baseline loss and step-by-step modify the loss function combination. Tab. 4 and Fig. 8 show that adding the AMSS and Bayer loss steadily improves visual quality and quantitative results. RepUNet with joint loss achieves the best results, outperforming the baseline by 0.5 dB in PSNR and by 0.16 in $\Delta E$ on average. As Fig. 8 shows, alignment-free and motion-aware short-exposure-first selection loss (AMSS-Loss) effectively suppresses the ghosting artifacts (see columns 3 and 4). Meanwhile, our proposed Bayer loss can alleviate the color cast (see columns 2 and 3).
@@
6 Conclusion
In the paper, we proposed a Topological Convolution Block (TCB) for efficient and light-weight HDR design for mobile devices. Based on the proposed TCB, we further designed RepUNet, aiming at balancing hardware efficiency and PSNR/SSIM indexes. Furthermore, We propose a novel computational photography based pipeline for raw HDR image formation and construct a real-world raw HDR dataset – RealRaw-HDR. Meanwhile, we designed plug-and-play alignment-free and motion-aware short-exposure-first selection loss efficient mitigate ghost artifacts. Our empirical evaluation validates the effectiveness of the proposed LDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.

### v6

@@ V-B Comparison with the Other Methods
Evaluation on HDR sensor dataset. To validate the validity of our method on the real-world HDR sensor dataset, we utilize the Chen test dataset [5] for cross-validation, which has raw images captured by a Sony IMX267 image sensor. Our method achieves comparable performance in visual quality and quantitative metrics compared to previous methods. The visual results from tests on the HDR sensor raw dataset (as shown in Fig. 6) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from color casts. Furthermore, results in Tab. III reveal that compared to AHDRNet [46], our method yields an improvement of more than 0.35 dB and 0.05 gain in PSNR and $\Delta E$, respectively, for scenes with an exposure ratio of 8. On average, our method attains gains exceeding 0.1 dB and 0.15 in PSNR and $\Delta E$. In addition, as shown in Tab. IV, RepUNet far outperforms the other methods in terms of PU-PSNR metrics, outperforming the suboptimal method by 1 dB. While our method yields a slightly lower score on ColorVideoVDP, it consistently outperforms competing methods on other HDR metrics (e.g., PSNR, SSIM, LPIPS), and offers lower computational cost and faster inference.
@@
TABLE V: Performance comparison of different HDR models on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images. The FLOPs are measured on the raw image of $7808\times 5824$ resolution. The best and second-best performances are in bold and underlined, respectively.
@@
Evaluation on FUJI raw dataset. We then evaluate our method on the FUJI raw datasets, which are real-world bracketed exposure raw images captured by the FUJI-FILM GFX50S II camera. Compared with previous methods, our method achieves state-of-the-art performance in visual quality and quantitative metrics. Fig. 7 compares results from HDR scenes, where our method achieves significantly better visualization. Our method can recover both fine details in overexposed regions and rich colors in underexposed areas without introducing artifacts (see rows 1 and 2). Compared to AHDRNet, our method can effectively remove noise and preserve the structure of dark regions. Tab. V shows that our method is far less computationally (FLOPs) intensive than AHDRNet [46] and HDR-Transformer at similar PSNR. Notably, the alignment module in DeepHDR, AHDRNet, HDR-Transformer, and NHDRRNet requires many line buffers, making it challenging to deploy on resource-limited edge devices. HDR-Transformer fails to perform inference even on RTX 3090 devices. In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images.
@@
V-C Ablation Study
This section investigates the raw SDR-HDR pair formation pipeline and the importance of different components in the whole RepUNet. We ablate the baseline model step by step and compare the performance differences.
Generalization of our SDR-HDR pair formation pipeline. Our raw SDR-HDR pair formation pipeline is proposed to generate paired raw SDR-HDR data, but can also be adapted to generate paired sRGB HDR data. To demonstrate such generalization, we transform the collected RealRaw-HDR dataset with a fixed ISP pipeline into the sRGB color space, named the Raw2RGB-HDR dataset. For comparison, we train the sRGB HDR method AHDRNet [46] on our Raw2RGB-HDR dataset and Kalantari dataset [18] (taking the first two exposures as input, 74 pairs of images), respectively. The test dataset is from the Kalantari dataset. Results in Tab. VI show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on the Kalantari dataset by 2.86 dB in PSNR. The performance gains benefit from an efficient and user-friendly data acquisition pipeline that generates more trainable data pairs. The results demonstrate that our data pipeline is also effective in generating paired SDR-HDR data in sRGB space.
TABLE VI: We train the HDR sRGB method AHDRNet on our Raw2RGB-HDR and Kalantari datasets, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR.
@@ V-C Ablation Study
TABLE VII: Quantitative comparisons of different loss functions. AMSS-Loss represents the alignment-free and motion-aware short-exposure-first selection loss.
[…]
@@
TABLE VIII: Reparameterization ablation results. The FLOPs and run times are measured on the raw image with a 4K resolution.
Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
@@
Loss functions. To test the effects of alignment-free and motion-aware short-exposure-first selection loss and Bayer loss, we set the $L2$ joint $L_{\textrm{ssim}}$ loss as the baseline and step-by-step modify the loss function combination. Tab. VII and Fig. 8 show that adding the AMSS and Bayer loss steadily improves visual quality and quantitative results. RepUNet with joint loss achieves the best results, outperforming the baseline by 0.5 dB in PSNR and by 0.16 in $\Delta E$ on average. As Fig. 8 shows, alignment-free and motion-aware short-exposure-first selection loss (AMSS-Loss) effectively suppresses the ghosting artifacts (see columns 3 and 4). Meanwhile, our proposed Bayer loss can alleviate the color cast (see columns 2 and 3).
[…]
VI Conclusion
In the paper, we proposed a Topological Convolution Block (TCB) for an efficient and lightweight HDR design that may be suitable for mobile devices. Based on the proposed TCB, we further designed RepUNet, aiming at balancing hardware efficiency and PSNR/SSIM indexes. Furthermore, we propose a novel computational photography-based pipeline for raw HDR image formation and construct a real-world raw HDR dataset – RealRaw-HDR. Meanwhile, we designed a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate ghost artifacts. Our empirical evaluation validates the effectiveness of the proposed SDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.
VII Limitations and Discussion
In this work, we propose an efficient and lightweight HDR network for dual-exposure HDR sensors. We have designed the TCB to prioritize computational efficiency, but real-world mobile deployments require additional hardware-specific adaptations. We will address this issue in future work through real-device testing. In addition, the proposed data synthesis pipeline only supports dual-exposure HDR sensor settings, and we will explore data synthesis pipelines for a wider range of triple-exposure HDR sensors in subsequent research. Additionally, we will focus on exploring more complex application scenarios such as flickering lights and large motion scenes.
Appendix
In this Appendix, we provide additional results and analysis.
VIII Supplementary Materials
Fig. 9 to Fig. 12 show additional visual comparisons of our method with state-of-the-art HDR reconstruction methods. As shown in these figures, DeepHDR, NHDRRNet, and SGN all have obvious ghosting and color residuals, especially for moving color objects, and the reconstruction results will have severe ghosting. Our proposed RepUNet can reconstruct HDR images without ghosting and color bias, and we can effectively remove the noise in the reconstructed HDR.
Figure 9: Visual comparisons with the state-of-the-art methods on the actual HDR sensor raw dataset from Chen’s dataset [5].
Figure 10: Visual comparison of state-of-the-art HDR reconstruction methods on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.
Figure 11: Visual comparison of state-of-the-art HDR reconstruction methods on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.
Figure 12: Visual comparison of state-of-the-art HDR reconstruction methods on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.